# HPLT 3.0: Very Large-Scale Multilingual Resources for LLM and MT Mono- and Bi-lingual Data, Multilingual Evaluation, and Pre-Trained Models

### v1

@@ Abstract
Laurie Burchell, Lucas Charpentier, Pinzhen Chen, Mariya Fedorova, Ona de Gibert,
Barry Haddow, Jan Hajič, Jindrič Helcl, Andrey Kutuzov, Zihao Li, Risto Luukkonen,
Bhavitvya Malik, Vladislav Mikhailov, Amanda Myntti, Dayyán O’Brien, Lucie Poláková,
Sampo Pyysalo, Gema Ramírez Sánchez, Janine Siewert, Pavel Stepachev,
Jörg Tiedemann, Teemu Vahtola, Fedor Vitiugin, Tea Vojtěchová, Jaume Zaragoza
@@
University of Turku, Department of Computing
@@ 1.  Introduction
Following the examples of FineWeb and HPLT 2.0, we make this resource publicly available under permissive licenses. [^1] https://hplt-project.org/datasets/v3.0 We further share the refined open-source processing pipeline, a novel multilingual evaluation framework, as well as various collections of language models pre-trained on HPLT 3.0 data. Finally, while our focus in this report is on the monolingual data and models, we also briefly report on ongoing work to derive novel bilingual datasets for 28 language pairs, provide associated machine translation models, and synthesize additional pre-training data for underrepresented languages by machine translation of very high-quality English documents. In our view, it is the totality of generally available and very large-scale resources and the documentation of the underlying processes that bears promise of “democratizing” the current LLM and MT landscape. In-depth technical and experimental details will be provided through appendices upon publication.
@@ 3.  Monolingual Data Preparation
Extraction of high-quality and richly annotated text from raw web archives proceeds through a sequence of refinement and filtering steps. Figure 1 shows the main components of our data preparation, which is a refined and extended version of the open-source HPLT pipeline by Burchell et al. (2025). The following paragraphs provide a high-level overview of the refinements we make to the original HPLT 2.0 pipeline.
@@ 7.  Multilingual LLM Evaluation
We develop HPLT 3.0-E, a framework for automated large-scale multilingual evaluation designed to systematically compare and refine data preparation choices across nine selected languages shown in footnote 4. These languages are chosen to ensure both availability of native speakers in our development team and a minimum level of diversity in terms of language resources, families, and scripts. Similar to Penedo et al. (2024, 2025), we pretrain separate language models per language using an otherwise fixed pretraining setup, and evaluate them at regular checkpoint intervals (every 1B tokens) in a 0-shot regime, carefully selecting tasks that meet the pretraining evaluation signal criteria below.
7.1.  HPLT 3.0-E: Framework
HPLT 3.0-E includes 127 language understanding and generation tasks, each supporting 3–7 human-written prompts. We aim to cover different task categories in all languages: entailment, commonsense reasoning, language-specific & world knowledge, paraphrasing, reading comprehension, sentiment analysis, toxicity detection, and truthfulness. HPLT 3.0-E integrates with LM Evaluation Harness Gao et al. (2024), for experimental flexibility and replicability.
@@
HPLT 3.0-E enables evaluation across 500+ prompts that have diverse structure and answer formatting to mitigate prompt sensitivity, a model limitation where variations in prompt formulation can affect downstream performance (e.g. Pezeshkpour and Hruschka, 2024; Sclar et al., 2024). We adapt single-prompt benchmarks (IberoBench, FrenchBench, and UkrainianBench) to the multi-prompt design through (i) manual translation of English prompts from PromptSource Bach et al. (2022) and (ii) prompt creation by native speakers.
@@ 7.2.  Results
In line with Penedo et al. (2025), we find that tasks for lesser-resourced languages, notably Basque and Galician, are unsuitable for pretraining evaluation due to potential difficulty and lack of monotonic performance progression during pretraining. We thus report our key findings on a final suite of 27 selected tasks across seven remaining languages.
@@
As another application of HPLT 3.0-E, Figure 3 shows the performance of three Spanish models trained on different subsets of HPLT 3.0 training data selected according to WDS levels; see § 3 above. In these experiments, five Spanish tasks were selected as informative for all models. Here, “random” sampling represents the default approach, drawing uniformly on the full dataset, while “top” and “bottom” take advantage of the sorting by WDS levels and sequentially draw 100B training tokens from either end of the dataset. Low WDS levels clearly lead to inferior model performance, while sampling from only the “top” does not cearly improve over the full dataset, possibly owing to overly limited diversity. We observe similar patterns in French, the other language supported by HPLT 3.0-E where there are hundreds of billions of available training data.
@@ 8.  Monolingual Encoder–Decoders
In this section, we apply the HPLT 3.0 dataset to train and evaluate 57 language-specific monolingual encoder–decoder language models, following the T5-base architecture Raffel et al. (2020).
@@
As regards the second purpose, we note the only available encoder–decoder with multilingual capabilities is mT5 pretrained on the mC4 corpus Xue et al. (2021), and its instruction-tuned derivatives, e.g. Muennighoff et al. (2023b); Üstün et al. (2024). Our collection of T5-like models is trained on more recent HPLT 3.0 datasets, which we believe to be substantially cleaner and more high-quality than mC4 (see our evaluation below). Thus, we hope that the NLP community will benefit from these novel models for the 57 languages that we currently support. In addition, language-specific models can be used to conduct comparative studies of LLM learning process (we also publish intermediate checkpoints). Their size ($\approx$275M parameters each) supports for easy and computationally cheap deployment. Our models follow the setup of Samuel et al. (2023).
Selection of Languages
@@
We also conducted additional evaluations on the English MultiBLIMP: The original monolingual T5-base performed better (93.5% accuracy), than the multilingual one; instruction-tuned derivatives mt0-xl (90.5% accuracy) and aya-101 (86.6% accuracy) further proved the finding by Jumelet et al. (2025) that fine-tuning worsens a model’s BLIMP performance; the most recent English T5-base model t5gemma-b-b-ul2 showed surprisingly low result (66.5% accuracy); however, its training objective was more sophisticated Tay et al. (2023) than in the original T5. We leave further adjusting our inputs to it for future work.
@@ 14.  Acknowledgments
We thank Étienne Simon (UiO) and Daryna Dementieva (TUM) for their contribution to our prompt collection for French and Ukrainian.

### v2

@@ Abstract
Laurie Burchell, Lucas Charpentier, Pinzhen Chen, Mariia Fedorova, Ona de Gibert,
Barry Haddow, Jan Hajič, Jindřich Helcl, Andrey Kutuzov, Veronika Laippala, Zihao Li,
Risto Luukkonen, Bhavitvya Malik, Vladislav Mikhailov, Amanda Myntti,
Dayyán O’Brien, Lucie Poláková, Sampo Pyysalo, Gema Ramírez Sánchez,
Janine Siewert, Pavel Stepachev, Jörg Tiedemann, Teemu Vahtola,
Dušan Variš, Fedor Vitiugin, Tea Vojtěchová, Jaume Zaragoza
@@
TurkuNLP, University of Turku, Department of Computing
@@ 1.  Introduction
Following the examples of FineWeb and HPLT 2.0, we make this resource publicly available under permissive licenses. [^1] https://hplt-project.org/datasets/v3.0 We further share the refined open-source processing pipeline, a novel multilingual evaluation framework, as well as various collections of language models pre-trained on HPLT 3.0 data. Finally, while our focus in this report is on the monolingual data and models, we also briefly report on ongoing work to derive novel bilingual datasets for 28 language pairs, provide associated machine translation models, and synthesize additional pre-training data for underrepresented languages by machine translation of very high-quality English documents. In our view, it is the totality of generally available and very large-scale resources and the documentation of the underlying processes that bears promise of “democratizing” the current LLM and MT landscape. In-depth technical and experimental details will be provided through appendices in a forthcoming extended version of this report.
@@ 3.  Monolingual Data Preparation
Extraction of high-quality and richly annotated text from raw web archives proceeds through a sequence of refinement and filtering steps. Figure 1 shows the main components of our data preparation, which is an updated and extended version of the open-source HPLT pipeline by Burchell et al. (2025). The following paragraphs provide a high-level overview of the refinements we make to the original HPLT 2.0 pipeline.
@@ 7.  Multilingual LLM Evaluation
We develop HPLT-E, a framework for automated large-scale multilingual evaluation designed to systematically compare and refine data preparation choices across nine selected languages shown in footnote 4. These languages are chosen to ensure both availability of native speakers in our development team and a minimum level of diversity in terms of language resources, families, and scripts. Similar to Penedo et al. (2024, 2025), we pretrain separate language models per language using an otherwise fixed pretraining setup, and evaluate them at regular checkpoint intervals (every 1B tokens) in a 0-shot regime, carefully selecting tasks that meet the pretraining evaluation signal criteria below.
7.1.  HPLT-E: Framework
HPLT-E includes 127 language understanding and generation tasks, each supporting 3–7 human-written prompts. We aim to cover different task categories in all languages: entailment, commonsense reasoning, language-specific & world knowledge, paraphrasing, reading comprehension, sentiment analysis, toxicity detection, and truthfulness. HPLT-E integrates with LM Evaluation Harness Gao et al. (2024), for experimental flexibility and replicability.
@@ 7.1.  HPLT-E: Framework
HPLT-E enables evaluation across 500+ prompts that have diverse structure and answer formatting to mitigate prompt sensitivity, a model limitation where variations in prompt formulation can affect downstream performance (e.g. Pezeshkpour and Hruschka, 2024; Sclar et al., 2024). We adapt single-prompt benchmarks (IberoBench, FrenchBench, and UkrainianBench) to the multi-prompt design through (i) manual translation of English prompts from PromptSource Bach et al. (2022) and (ii) prompt creation by native speakers.
@@ 7.2.  Results
In line with Penedo et al. (2025), we find that tasks for lesser-resourced languages, notably Basque and Galician, are unsuitable for pretraining evaluation due to potential difficulty and lack of monotonic performance progression during pretraining. We thus report our key findings on a final suite of 26 selected tasks across seven remaining languages.
@@
As another application of HPLT-E, Figure 3 shows the performance of three Spanish models trained on different subsets of HPLT 3.0 training data selected according to WDS levels; see § 3 above. In these experiments, five Spanish tasks were selected as informative for all models. Here, “random” sampling represents the default approach, drawing uniformly on the full dataset, while “top” and “bottom” take advantage of the sorting by WDS levels and sequentially draw 100B training tokens from either end of the dataset. Low WDS levels clearly lead to inferior model performance, while sampling from only the “top” does not clearly improve over the full dataset, possibly owing to overly limited diversity. We observe similar patterns in French, the other language supported by HPLT-E where there are hundreds of billions of available training data.
@@ 8.  Monolingual Encoder–Decoders
In this section, we apply the HPLT 3.0 dataset to train and evaluate 57 language-specific monolingual encoder–decoder language models, following the T5-base architecture Raffel et al. (2020). This novel family of models, including intermediate checkpoints, is publicly available. [^8] https://huggingface.co/collections/HPLT/hplt-30-t5-models
@@
As regards the second purpose, we note the only available encoder–decoder with multilingual capabilities is mT5 pretrained on the mC4 corpus Xue et al. (2021), and its instruction-tuned derivatives, e.g. Muennighoff et al. (2023b); Üstün et al. (2024). Our collection of T5-like models is trained on more recent HPLT 3.0 datasets, which we believe to be substantially cleaner and more high-quality than mC4 (see our evaluation below). Thus, we hope that the NLP community will benefit from these novel models for the 57 languages that we currently support. In addition, language-specific models can be used to conduct comparative studies of LLM the learning process (we also publish intermediate checkpoints). Their size ($\approx$275M parameters each) allows for easy and computationally cheap deployment. Our models follow the setup of Samuel et al. (2023).
Selection of Languages
@@
We also conducted additional evaluations on the English MultiBLIMP: The original monolingual T5-base performed better (93.5% accuracy), than the multilingual one; instruction-tuned derivatives mt0-xl (90.5% accuracy) and aya-101 (86.6% accuracy) further proved the finding by Jumelet et al. (2025) that fine-tuning tends to worsen BLIMP performance; the most recent English T5-base model, t5gemma-b-b-ul2, showed surprisingly low result (66.5% accuracy); however, its training objective was more sophisticated Tay et al. (2023) than in the original T5. We leave further adjusting our inputs to it for future work.
@@ 14.  Acknowledgments
We thank Étienne Simon (UiO) and Daryna Dementieva (TUM) for their contribution to our prompt collection for French and Ukrainian and Erik Henriksson, Erofili Psaltaki, and Otto Tarkka (all UTU) for their contributions to the manual inspection.